OpenStudy (anonymous):

A right triangle has one side of length 8 and a hypotenuse of length 10. How long is the third side of the triangle?

OpenStudy (anonymous):

\[Other \; Side = \sqrt{10^2 - 8^2} = \sqrt{100 - 64} = \sqrt{36} = 6\]

OpenStudy (anonymous):

\[\huge \color{blue}{Hypotenuse^2 = Base^2 + Height^2}\]

OpenStudy (anonymous):